Suppose that the firm operates in a perfectly competitive market. The market price of his

product is$10. The firm estimates its cost of production with the following cost function:

TC=10q-4q2+q3
1
Expert's answer
2021-09-26T19:20:55-0400

From the above information, we can determine the level of output that this firm should produce to maximize its profit.

Under perfect competition MR = P.

where P = MR = MC.

P = $10

MC = derivative of the Total Cost (TC) in relative to quantity

TC "=10q-4q^2+q^3"

MC "=" "\\frac{\\partial TC} {\\partial q} \n\u200b\t\n = 10 \u2013 8q + 3q2"

MC "=10-8q+3q^2"

to derive the level of output that maximizes the firm’s profit,

P = MC

"10=10-8q+3q^2"

"3q^2-8q=0"

"3q^2=8q"

"3q=8"

"q=8\/3"

"=2.667"

The level of output that the firm should produce to maximize its profit: q = 2.667